OpenStudy (anonymous):

the perimeter is 90 in. and the width is 19 in.

OpenStudy (anonymous):

perimeter of a what? welcome to openstudy!

OpenStudy (anonymous):

rectangle i have to find the unknown measurement

OpenStudy (anonymous):

Solve the following for L: 2 L + 2*19 = 90

OpenStudy (anonymous):

what to do after take the two in the middle & multiply ?

OpenStudy (anonymous):

2L + 38 = 90 2L = 90-38

OpenStudy (anonymous):

next ? you take it from both sides dont you ?

OpenStudy (anonymous):

Divide each side by 2. The LHS is then L and the RHS is 26.

OpenStudy (anonymous):

thanks ! you helped me lots :)

OpenStudy (anonymous):

You are welcome and thank you for the medal.